# AdaNCA: Neural Cellular Automata As Adaptors For More Robust Vision Transformer

### v4

@@ 5 Limitation
AdaNCA has certain limitations. First, the AdaNCA-equipped ViTs cannot adapt to unseen recurrent steps of AdaNCA, limiting the generalization ability. For example, if the training step range of AdaNCA is [3,5], it cannot produce meaningful results with the test step being 6. Moreover, AdaNCA introduces non-neglectable computation into the original architecture. Furthermore, our experiments are conducted on ImageNet1K with the image size being 224 $\times$ 224. Whether AdaNCA can lead to impressive improvement on larger-scale problems, e.g., ImageNet22K, remains a question. The size of the images can also affect the efficiency of token interaction.
@@ 6 Broader Impact
AdaNCA contributes to more robust ViTs, facilitating their usage in real-world scenarios. Moreover, we bridge two powerful models, NCA and ViT, on large-scale image classification, potentially encouraging research dedicated to their synergistic combination in more practical settings. Our finding on AdaNCA improving network redundancy can stimulate more works on architectural robustness in deep learning, which involves increasing the redundancy to enhance robustness.
@@ 7 Conclusion
We propose AdaNCA, an efficient Neural Cellular Automata (NCA) that, when inserted between their middle layers, improves ViT performances and robustness against adversarial attacks as well as out-of-distribution inputs. We motivate our design by connecting NCA and ViT in terms of token interaction modeling, and propose Dynamic Interaction to improve the computational efficiency of standard NCA. Exploiting the training strategies and design choices in NCA, i.e., stochastic update, random steps, and multi-scale interaction, we further improve the AdaNCA-enhanced ViTs’ clean accuracy and robustness. To decide the placement of AdaNCA, we propose the Set Cohesion Index that quantifies the network redundancy via layer similarity and conclude that AdaNCA should be inserted between two layer sets that consist of redundant layers. Our results demonstrate that AdaNCA consistently improves ViTs performances and robustness. Evidence suggests that the mechanism by which we obtain improvement reduces the sensitivity of ViTs to certain types of noise and makes the noise-resilient ability of ViTs closer to humans.
@@ C.10 Visualization of attention maps
C.11 Additional results on same model architecture but with more layers ($\star$ models)
@@
C.12 Additional results on layer similarity structure
@@
C.13 Additional results on category-wise mCE on ImageNet-C
@@
C.14 Additional results on noise sensitivity examination
@@
C.15 Datasets information and license
@@
C.16 Model and code license
@@
Appendix D Analysis on the efficiency improvement of AdaNCA
In our experiments, we do not conduct experiments on comparing the concatenation scheme and our weighted sum scheme in the Interaction stage. It is because the concatenation scheme will result in a huge increase in the number of parameters as well as FLOPs, rendering comparison unfair and making the source of performance improvement undecidable. Concretely, with both the token and hidden dimensions of the MLP set at $C$, the parameter count and FLOPs required to update a single token are nearly $5C^{2}$. For a standard base-level ViT configuration with $C=768$ and token number $N=196$, this results in near $0.6G$ FLOPs per iteration for a single NCA, and incorporating recurrence further increases the computational demands. If we adopt the concatenation scheme in the RVT-B-AdaNCA model, it will result in $\sim$10M more parameters and $\sim$4G more FLOPs during test. Hence, we decide not to include the comparison between the original Interaction scheme and our Dynamic Interaction. However, we do conduct an ablation in Section 4.3 on mean aggregation versus our weight sum scheme and demonstrate the effectiveness of our proposed Dynamic Interaction in improving model performance and robustness.

### v5

@@ 5 Limitation
AdaNCA has certain limitations. First, the AdaNCA-equipped ViTs cannot adapt to unseen recurrent steps of AdaNCA, which limits the generalization ability. For example, if the training step range of AdaNCA is [3,5], it cannot produce meaningful results with the test step being 6. AdaNCA introduces a non-negligible computation into the original architecture. Our experiments are conducted on ImageNet1K with the image size of 224 $\times$ 224. Whether AdaNCA can lead to impressive improvements on larger-scale problems, e.g., ImageNet22K, remains a question. The size of the images can also affect the efficiency of token interaction.
@@ 6 Broader Impact
AdaNCA contributes to more robust ViTs, facilitating their usage in real-world scenarios. We bridge two powerful models, NCA and ViT, on large-scale image classification, potentially encouraging research dedicated to their synergistic combination in more practical settings. Our findings on AdaNCA improving network redundancy can stimulate more works on architectural robustness in deep learning that involves increasing the redundancy to enhance robustness. In the context of this paper, we believe that AdaNCA does not introduce any significant negative implications.
@@ 7 Conclusion
We have proposed AdaNCA, an efficient Neural Cellular Automata (NCA) that, when inserted between their middle layers, improves ViT performances and robustness against adversarial attacks as well as out-of-distribution inputs. We design our model by connecting NCA and ViT, in terms of token interaction modeling, and we propose Dynamic Interaction to improve the computational efficiency of standard NCA. Exploiting the training strategies and design choices in NCA, i.e., stochastic update, random steps, and multi-scale interaction, we further improve the AdaNCA-enhanced ViTs’ clean accuracy and robustness. To decide the placement of AdaNCA, we propose the Set Cohesion Index that quantifies the network redundancy via layer similarity and conclude that AdaNCA should be inserted between two layer sets that consist of redundant layers. Our results demonstrate that AdaNCA consistently improves ViTs performances and robustness. Evidence suggests that the mechanism by which we obtain improvement reduces the sensitivity of ViTs to certain types of noise and makes the noise-resilient ability of ViTs similar to that of humans.
@@ C.1 AdaNCA settings for each model
C.1.1 Discussion on the choice of AdaNCA steps
Our choice of AdaNCA steps differs from all previous NCA models, which typically iterate for hundreds of steps. We make this compromise as the computational costs increase linearly as the step grows. We aim to minimize the increase in the number of parameters and FLOPS for scalability and we do not want the source of improvement to merely stem from the increase in the size and computation of the models. We show that more NCA steps will indeed contribute to the model’s performance in Table 7, while it introduces extra FLOPS. The model is the same as in Section 4.3 which is trained using a range of steps being $[3,5]$. We want to underscore that it is the architecture and evolution scheme that defines an NCA as introduced in Section 3.1, instead of the number of its recurrent steps. Admittedly, fewer steps will lead to a coarser path to the target state, and can potentially undermine the model’s performance. Notably, with our scheme, AdaNCA has achieved generally strong robustness improvements compared to the baselines. It indicates that our choice of the recurrent steps of AdaNCA is a good balance between computational costs and performance.
Table 7: Increasing the number of recurrent steps of AdaNCA during testing can contribute to the model performance.
# Params (M) FLOPS (G) Clean Acc. ($\uparrow$) Attack Failure Rate ($\uparrow$)
Step=4 27.94 4.7 87.18 22.35
Step=5 27.94 4.8 87.22 23.46
@@ C.10 Visualization of attention maps
C.11 Integrating AdaNCA into pre-trained ViT models
In our experiments, we train all models from scratch. Implementing AdaNCA as a plug-and-play module for pre-trained ViT models would certainly improve the training speed. To explore this, we experiment by inserting AdaNCA into a pre-trained Swin-base model on ImageNet1K by 1) freezing all ViT layers; 2) training only the boundary layers, where boundary layers are the layers before and after the insert position of AdaNCA; 3) Finetuning all layers. Results are given in Table 14. None of the schemes perform as well as training from scratch. The results indicate that the current NCA is able to adapt to such a scheme. However, it may struggle to effectively transmit information between two pre-trained ViT layers, as these layers have already established strong connections. In contrast, training the model from scratch allows NCA and ViT to synergistically adapt to feature variability, resulting in better overall performance. However, it is worth exploring in the future since the fine-tuning scheme can contribute to the performance.
Table 14: Integrating AdaNCA with a pre-trained ViT on ImageNet1K. All integration schemes perform worse than training Swin-B-AdaNCA from scratch.
Clean Acc. ($\uparrow$) IM-A ($\uparrow$) APGD-DLR ($\uparrow$) IM-R ($\uparrow$) IM-SK ($\uparrow$)
Swin-B 83.4 35.8 15.6 46.6 32.4
Swin-B-AdaNCA 83.7 36.0 25.1 48.2 35.5
Freeze All 83.1 30.9 11.5 46.2 32.5
Boundary Training 83.1 32.8 19.5 47.1 33.9
Finetune All 83.3 34.9 19.9 46.9 33.7
C.12 Additional results on comparison with current methods
We have presented the comparison between our AdaNCA and the current SOTA method, TAPADL models [23], in Section 4.1. Here, we provide an additional comparison with the ViTCA model [62]. In ViTCA, tokens interact with each other through local self-attention. We aim to compare our proposed Dynamic Interaction module with the interaction learning scheme in ViTCA. Results are given in Table 15. Despite having more parameters, the ViTCA-like scheme performs worse in clean accuracy and is on-par with AdaNCA in robustness. This indicates that it is promising to explore the possibility of incorporating ViTCA into our framework.
Table 15: Comparison with ViTCA. ViTCA performs worse in clean accuracy and is on-par with AdaNCA in robustness. Note that ViTCA has more parameters than our scheme Dynamic Interaction. The training setting is the same as in Section 4.3 in the main paper.
# Params (M) FLOPS (G) Clean Acc.($\uparrow$) Attack Failure Rate ($\uparrow$)
Dynamic Interaction 27.94 4.7 87.18 22.35
ViTCA, scale=1 28.48 4.8 86.78 19.47
ViTCA, scale=2 29.08 4.9 86.72 22.28
C.13 Additional results on same model architecture but with more layers ($\star$ models)
@@
C.14 Additional results on layer similarity structure
@@
C.15 Additional results on category-wise mCE on ImageNet-C
@@
C.16 Additional results on noise sensitivity examination
@@
C.17 Datasets information and license
@@
C.18 Model and code license
@@
Appendix D Extended ablation studies
In this section, we discuss the design of our AdaNCA and provide additional ablation studies on the designs of AdaNCA.
D.1 The dropout-like strategy
In Section 3.1.1, we introduce our approach of using the dropout-like strategy to perform the stochastic update during training and testing. While previous NCA works claim that the stochasticity can be preserved [44, 46] or can be simply switched off [62] during testing, the proposed dropout-like scheme is necessary in our case. First, We highlight that stochasticity during testing can hinder the evaluation of adversarial robustness by producing obfuscated gradients [3], leading to the circumvention of adversarial attacks. Table 19 illustrates this issue, where we test the classification accuracy under CW attack. Stochasticity also results in inconsistent outputs. This is problematic for practical image classification tasks where reliable output is critical, unlike applications focusing on visual effects [46, 62] or collective behaviors [55]. The change in clean accuracy in Table 19 indicates that given the same image, the stochasticity can lead to different decisions, hindering the deployment of the trained models in real-world scenarios. To further demonstrate the necessity of our scheme, we remove the compensation of the output magnitude during training, i.e., we remove the $\frac{1}{p}$ in Equation 7 during training and directly test the model with a synchronized update. The results are given in Table 20. The drop in clean accuracy and robustness suggests that downstream ViT layers struggle with varying NCA output magnitudes, indicating the usefulness of our proposed dropout-like method.
Table 19: Stochasticity during testing will give a false sense of adversarial robustness. We conduct the same experiments twice, which result in Trials 1 and 2. The training setting is the same as in Section 4.3 in the main paper.
Clean Acc. ($\uparrow$) CW Acc. ($\uparrow$)
Test-time Synchronous 87.18 9.48
Test-time Stochasticity, Trial 1 87.10 10.72
Test-time Stochasticity, Trial 2 87.06 10.60
Table 20: Removing $\frac{1}{p}$ in Equation 7 during training and directly testing with synchronized update. Our dropout-like compensation scheme improves AdaNCA’s performance.
Clean Acc. ($\uparrow$) Attack Failure Rate ($\uparrow$)
Original 87.18 22.35
No $\frac{1}{p}$ 86.84 21.94
D.2 Dynamic Interaction
Our motivation for developing the Dynamic Interaction module is the high dimensionality of feature vectors in modern ViT models. We underscore that any operations involving linear transformations of the concatenated interaction results will lead to drastic increases in computational costs, as shown in Table 21. Note that the concatenation scheme nearly doubles the FLOPS for RVT compared to the baseline, leading to difficulties in training. Such an increase renders scalability a challenge. Admittedly, more parameters can contribute to better performance, as shown in Table 22. Note, however, that our Dynamic Interaction achieves  70% of the improvements of the original NCA concatenation scheme (10.06/14.62) in robustness improvement and  60% improvement in the clean accuracy (0.62/1.06), with merely  10% of the parameters and FLOPS (0.35/2.99 for # Params and 0.2/2.3 for FLOPS). Therefore, our Dynamic Interaction scheme provides a good trade-off between the performance and computational costs. It is capable of scaling up, allowing us to insert AdaNCA into even larger models, such as current vision-language models where the token dimensionality is even higher. Moreover, we qualitatively showcase in Figure 12, that Dynamic Interaction helps robustify AdaNCA when facing noisy inputs.
Table 21: The concatenation scheme of the original NCA results in a large increase in the computational costs when the token dimensionality is high, leaving scaling up to large datasets and model sizes a challenge.
# Params (M) FLOPS (G)
Swin-B 87.8 15.4
Swin-B-AdaNCA 90.7 16.3
Swin-B-AdaNCA, Concat 115.5 24.7
RVT-B 88.5 17.7
RVT-B-AdaNCA 91.0 19.0
RVT-B-AdaNCA, Concat 112.3 34.0
Table 22: Performance of using our Dynamic Interaction and the concatenation scheme of original NCA on ImageNet100. Numbers in parentheses indicate performance improvement compared to the Baseline. Our method achieves a large portion of the performance improvement (60% 70%) brought by the concatenation scheme with much fewer computational costs ( 10%).
# Params (M) FLOPS (G) Clean Acc.($\uparrow$) Attack Failure Rate ($\uparrow$)
Baseline 27.59 4.5 86.56 12.29
Dynamic Interaction 27.94 (+0.35) 4.7 (+0.2) 87.18 (+0.62) 22.35 (+10.06)
Concat 30.58 (+2.99) 6.8 (+2.3) 87.62 (+1.06) 26.91 (+14.62)
Figure 12: Visualization of the developing token maps of AdaNCA using PCA. Without Dynamic Interaction, AdaNCA cannot recover from the noisy adversarial inputs. Equipped with it, the model can stick to the evolution path similar to the clean inputs.
D.3 Multi-scale Dynamic Interaction
Our motivation for developing multi-scale interaction is to perform more efficient token interaction learning over local scales since ViT already contains global information. Enlarging the neighborhood size will: 1) Complicate the process of selecting the neighbors to interact with; 2) Introduce excessive noise, making it difficult for tokens to accurately acquire neighbor information. 3) Repeatedly acquire the global information provided by ViT. The recurrence further amplifies the noise. In our ablation study in Section 4.3, we notice that increasing the number of scales to 3 will undermine the performance. Such an issue is also observed in previous work [62] where local self-attention is used for interaction learning. While in theory both Dynamic Interaction and self-attention can discount far-away information as the tokens can decide their own interaction neighborhood, overly large neighborhood size can lead to the problems mentioned above. Note that our model gains performance when $\mathcal{S}=2$ ($5\times 5$ neighborhood), indicating the usefulness of our multi-scale module. We further explore whether the multi-scale issue can be solved by increasing the model capacity. Specifically, we develop two additional schemes. The first is to replace the dilation in Dynamic Interaction by simply using larger filters. The other one enlarges the receptive field of the weight computation network $\mathcal{W}_{M}$ in Equation 9. Instead of using two $3\times 3$ convolution layers, we change the first layer to be a $5\times 5$ convolution. As a result, the receptive field of $\mathcal{W}_{M}$ becomes $7\times 7$, matching the one when $\mathcal{S}=3$. We use the same training setting as in Section 4.3. Results are given in Table 23. Although increasing model capacity might alleviate the noise issues, AdaNCA struggles with overly large neighborhoods.
Table 23: Overly large neighborhood size during token interaction will undermine the model performance as it complicates the process of figuring out which neighbors to interact with and adds too much noise during the process. While increasing the model capacity can alleviate such an issue, AdaNCA still struggles to handle too-noisy information when the scale is too high. Note that the larger filter scheme achieves on-par results with our method when the scale is 2, indicating that it does not provide additional useful information for token interaction at this scale.
# Params (M) FLOPS (G) Clean Acc.($\uparrow$) Attack Failure Rate ($\uparrow$)
Original, scale=3 27.94 4.7 86.60 20.44
Larger filter,scale=3 28.04 4.7 86.74 21.13
Larger $\mathcal{W}_{M}$ receptive field,scale=3 27.98 4.7 86.66 20.32
Original, scale=2 27.94 4.7 87.18 22.35
Larger filter,scale=2 27.96 4.7 87.14 22.31
Appendix E Differences between traditional NCA and AdaNCA
While AdaNCA is inspired by NCA, it does not fully exploit the designs of traditional NCA. We discuss three key differences between AdaNCA and NCA below:
  1. 1.
    Different initial states. NCA in image generation starts from either constant or randomly initialized cell states, typically referred to as the seed states, while AdaNCA receives the outputs from previous ViT layers. Hence, AdaNCA handles structured inputs at the very beginning.
  2. 2.
    Different usage of cell states. Traditional NCA does not use all cell states for accomplishing the downstream tasks. After certain steps of evolution, a subset of the cell states is extracted to perform the given task. The unused cell states, termed hidden states, facilitate cell communications as they can be used to store additional information [44]. The dimensionality of the hidden states is typically several times that of the input. When the cell dimensionality is high, as is the case in ViT, adding too many hidden states will bring too much computational costs. Moreover, cells can store effective enough information when they have high dimensionality. Therefore, we do not adopt the hidden states design in AdaNCA.
  3. 3.
    No pooling strategy. A critical component of traditional NCA is the pooling strategy. Instead of starting from the seed states in every epoch, NCA fetches the starting states from a pool, where the output states from previous epochs are stored. Through this strategy, NCA can explore much longer time steps without suffering from gradients or memory issues. While it is tempting to incorporate the pooling trick into our method, two major concerns hinder its practical implementation: 1) The previous NCA models start from their own outputs in the pooling strategy. In other words, the cell states in the pool are generated by the model itself. However, AdaNCA starts from the outputs of a ViT layer. Such a difference renders the pool trick hard to implement. 2) Taking a step back, even if there is a well-designed pooling strategy for AdaNCA, it will bring too much computational costs during testing. A single step of AdaNCA introduces non-trivial FLOPS during testing, as shown in Table 1. The ultimate goal of the pooling strategy is to ensure the stability of NCA in large time steps, while large time steps will slow down the inference.